What are the a, b and c terms in the quadratic equation with the answers -3 and -5?
Answer by stanbon(75887)

What are the a, b and c terms in the quadratic equation with the answers -3 and -5?
-------
f(x) = (x+3)(x+5) = x^2 +8x + 15
----------------
a = 1 ; b = 8 ; c = 15
--------------------
Cheers,
Stan H.